OpenStudy (anonymous):

A company builds computers. It costs $6,700 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x?

OpenStudy (nurali):

Given the information, you have two points on your line: (10, 6700) and (20, 12200) where the x-coordinate is computers and the y-coordinate is money! Now lets find a line that crosses these two points! First, lets find the slope: (12200-6700)/(20-10) = 550 Then, to find the y-intercept, you plug one of those points into y=mx+b with the slope that you just found. So: y=mx+b 6700 = 550*10 +b b = 6700 - 5500 = 1200 Now you have your slope and y-intercept, the two things you need to find your line! C(x) = 550x + 1200

OpenStudy (anonymous):

thank you and oh yeah

OpenStudy (anonymous):

@Kamizamurai

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Welcome to Openstudy.